OpenStudy (anonymous):

What is 5^5x-3 = 31?

OpenStudy (anonymous):

Did you write this correctly? Is it exponent 5 or 5x?

OpenStudy (anonymous):

Yes 5x-3, Thanks Luis!!

OpenStudy (anonymous):

Can you simplify further?

OpenStudy (anonymous):

Is this what the problem should look like 5(5x-3)=31? Divide both sides by 5 5x-3=31/5 Add 3 to both sides 5x=46/5 Divide both sides by 5 x= 46/25 or approx 1.8400000000000000000000000